OpenStudy (anonymous):

What is the approximate angle between two position vectors if their terminal points are (5, -2) and (7, 3)?

OpenStudy (anonymous):

Are you familiar with a "dot product?"

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

I know it but dont quite understand it

OpenStudy (anonymous):

A dot product has two very convenient methods of calculating it's value. Given the vectors A=(a,b) and B=(c,d): \[A . B =|A||B|Cos(\theta)\] and \[A.B = (a,b).(c,d) = ac+bd\] We can equate the two methods: \[ac+bd = |A||B|Cos(\theta)\] And then solve for the angle, theta.

OpenStudy (anonymous):

so like (5)(7)+(-2)(3)=cos theta?

OpenStudy (anonymous):

Got it. Thanks!

OpenStudy (anonymous):

Close, you dropped the |A| and |B| These are the "magnitudes" of A and B. It's just how long they are. You can use the Pythagorean theorem to find their magnitudes: If A = (a,b) \[|A| = \sqrt{a^2+b^2}\]